OpenStudy (anonymous):

Solve each exponential equation. Give the exact value for x. 9x = 1/81

OpenStudy (anonymous):

write 1/81 as 9^(-2) so you have 9^x = 9^(-2) Now equate exponents.

OpenStudy (charlotte123):

x = 1/729 that is one of the biggest denominators I seen in my life O_O

OpenStudy (charlotte123):

http://www.wolframalpha.com/input/?i=9x+%3D+1%2F81

OpenStudy (mathstudent55):

@lucascoco08 Did you mean 9^x = 1/81? If so, follow @Easyaspi314 If you meant 9x = 1/81, then follow @charlotte123 Since you state "Solve each exponential equation.", I assume you meant 9^x = 1/81

OpenStudy (charlotte123):

@mathstudent55 thank u!

OpenStudy (mathstudent55):

wlcm